fac = 1;
for i in range(1, 3):
fac *= i
print(fac)

(Python)

please explain why the answer is two and how to change it to other values like 4 or 5

please be complete and act as if you were explaining it to a small lost child

ty
Answer #1
This code computes product of numbers from 1 to n
Where n is the second argument of range - 1

This code computes the factorial of 2
So, answer is 1*2 = 2

if we change the number from 3 to 4
the it computes 1*2*3 = 6

if we change the number from 3 to 5
the it computes 1*2*3*4 = 24

if we change the number from 3 to 6
the it computes 1*2*3*4*5 = 120

the factorials in sequence are 1, 2, 6, 24, 120, .... 
There is no way to get 4 or 5 in the sequence